What is the slope of the line passing through the points A(5,-1) and B(-8,3)?

Answer by RAY100(1637)

Remember,,,,m=(y2-y1) / (x2-x1)
.
(5,-1),,,(-8,3)
.
m= { (3)-(-1)} / {(-8)-(5)} = 4/-13=-4/13
.
a simple sketch might help also
.